Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability question

Expert replies
by coscha » Wed Feb 03, 2010 11:59 am
If the probability that it will rain on any day in September is 0.3, caluclate the probability that in a given ween in September, it will rain on exactly three consecutive days.

I got 0.027...but the answer is supposed to be 0.0324

Help plz
Join the discussion
Source: — Problem Solving |

by ace_gre » Wed Feb 03, 2010 12:50 pm
Hi, Here is my approach.

Lets create a table with the days of the week. Consider a week from Sun - Sat. Days that rain are represented by R and no rain be X
There are 5 ways to rain on three consecutive days in a week.

Su M T W Th F Sa
R R R X X X X
X R R R X X X
X X R R R X X
X X X R R R X
X X X X R R R

Since required probability is exactly three rain days and 4 non rain days => 5 * 0.3 *0.3 *0.3 * (0.7)^4
==>0.0324
Join the discussion

by thephoenix » Wed Feb 03, 2010 12:59 pm
let us take this problem as if we are asked to place 3 R (rain days) and 4 Nr(non rain days) in 7 place

condition all R has to be togather

__1 __2 __3 __4 ___5 __6__7

now R can take any value from 1 to 5 ( in order to be consecutive)
it can be done in 5C1 ways=5
now 4Nr can be arranged in 4 left places in one way ( as order does not matters)

tot ways=5
now probabilty of a rain=.3
hence for 3 rain days P=.3*.3*.3
and for 4 Nr P=.7*.7*.7*.7

fnal P=5*(.3^)*(.7^4)=.0324
Join the discussion

by ajith » Wed Feb 03, 2010 1:00 pm
coscha wrote:If the probability that it will rain on any day in September is 0.3, caluclate the probability that in a given week in September, it will rain on exactly three consecutive days.

I got 0.027...but the answer is supposed to be 0.0324

Help plz
The rain can start with first day (1st day 2nd day and 3rd day rains) Second day , third day or fourth day and 5th day

In case the rain starts with first day it has to rain first day second and third day, additionally it should not rain on 4th 5th 6th and 7th days ; probability = 1/3*1/3*1/3*2/3*2/3*2/3*2/3 = 16 /3^6

similarly probability for rains starting on second day is = 16/3^6... and so on so forth till 5th day

total probability = 80/3^6

Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by ace_gre » Wed Feb 03, 2010 3:20 pm
Hey ajith, 0.3 = 3/10 and not 1/3..and probability of not raining is 7/10 not 2/3 ? I guess you are tired :)
Join the discussion

by ajith » Wed Feb 03, 2010 9:13 pm
ace_gre wrote:Hey ajith, 0.3 = 3/10 and not 1/3..and probability of not raining is 7/10 not 2/3 ? I guess you are tired :)
Ahhh I get it... I guess i am tired too, Thanks
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by coscha » Thu Feb 04, 2010 8:50 am
Thanks alot guys
Join the discussion